OpenStudy (anonymous):

if1/x,1/y,1/z are in A.P then show that xy.zx,yz are in A.P pls show me the steps of the sum

terenzreignz (terenzreignz):

Well, what does it mean for these three to be in arithmetic progression?

terenzreignz (terenzreignz):

These three, to be precise: \[\Large \frac1x \qquad \frac1y \qquad \frac1z\]

OpenStudy (anonymous):

pls read the question once more

terenzreignz (terenzreignz):

What can you immediately conclude if you know that those three ^ are in an arithmetic progression?

OpenStudy (anonymous):

can u prove wht i have asked for

terenzreignz (terenzreignz):

Maybe... I don't really know... I haven't gone round to actually doing it yet... I'd like you to do it with me, so that we'll learn together, though :)

terenzreignz (terenzreignz):

Come one, what does it mean, for things to be in an arithmetic progression (AP)?

OpenStudy (anonymous):

i have o time to explain it to right now

terenzreignz (terenzreignz):

Well, just a rough definition of what makes an AP will do...

OpenStudy (anonymous):

which class are u in??

terenzreignz (terenzreignz):

LOL If these three \[\Large \frac1x \qquad \frac1y \qquad \frac1z\] are in arithmetic progression, then \[\Large \frac1y - \frac1x = \frac1z - \frac1y\] correct?

OpenStudy (anonymous):

yap

terenzreignz (terenzreignz):

Well then, what I want you to do is simplify both sides of \[\Large \frac1y - \frac1x = \frac1z - \frac1y\] so that there is only one fraction on each side...

OpenStudy (anonymous):

ok.wait

OpenStudy (anonymous):

its not working

terenzreignz (terenzreignz):

What's not working?

OpenStudy (anonymous):

wht u asked me to do

terenzreignz (terenzreignz):

Oh you of little faith :P I've already done it ^_^ Now all that's left is for you to be able to do it too :)

terenzreignz (terenzreignz):

Come on, just trust me ;)

OpenStudy (anonymous):

ok.wait

OpenStudy (anonymous):

ok i have got y-x/xy=y-z/yz after simplifying it

terenzreignz (terenzreignz):

Nope... Your left side is off... once you realise that \[\Large \color{red}{\frac{y-x}{xy}=\frac1x - \frac1y}\color{blue}{\neq \frac{1}y-\frac1x}\]

terenzreignz (terenzreignz):

A rather small error, but significant, nevertheless... you got it backwards... \[\Large \frac{x-y}{xy}=\frac{y-z}{yz}\] ok? :P

OpenStudy (anonymous):

after that wht should i do

terenzreignz (terenzreignz):

Now, it'd be nice if both sides had the same denominator... a good candidate would be the denominator xyz. On the left side, we may multiply the fraction \(\Large \frac{z}z\) without any worry, since this is effectively just 1. On the right side, we may multiply the fraction \(\Large \frac{x}{x}\) with similar results... \[\Large \frac{z}{z}\cdot\frac{x-y}{xy}=\frac{x}{x }\cdot\frac{y-z}{yz}\]

terenzreignz (terenzreignz):

Please distribute :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

its zx-zy /zxy=xy-xz/xyz

terenzreignz (terenzreignz):

mhmm :) Let me just rewrite that for you ^_^ \[\Large \frac{zx-yz}{xyz}=\frac{xy-zx}{xyz}\] Notice that the two sides now have the same denominator, and so they may be cancelled... \[\Large \frac{zx-yz}{\cancel{xyz}}=\frac{xy-zx}{\cancel{xyz}}\] \[\Large zx - yz = xy - zx\]

OpenStudy (anonymous):

then...........

terenzreignz (terenzreignz):

then......? If you like, you can multiply -1 to both sides (actually, it's unnecessary, but meh) \[\Large yz - zx = zx - xy\]

OpenStudy (anonymous):

its not done ,dude

terenzreignz (terenzreignz):

And why not?

OpenStudy (anonymous):

so from this how can u prove that xy.zx,yz are in A.P

terenzreignz (terenzreignz):

What is the definition of an AP? That's why that's the first thing I asked you... so that you're clear on your objectives... the numbers a, b, and c form an arithmetic progression if b - a = c - b

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

So... We have now shown that yz - zx = zx - xy What don't you get now?

OpenStudy (anonymous):

yes,lol i got it .hey thank u so much dude

terenzreignz (terenzreignz):

no problem